OpenStudy (anonymous):

Find the extreme values of f(x)=Sqx+3

OpenStudy (ranga):

\[f(x) = \sqrt{x} ~~or~~ f(x) = \sqrt{x + 3}\]

OpenStudy (anonymous):

the second one

OpenStudy (ranga):

It is square root and not a square, right?

OpenStudy (anonymous):

\[\sqrt{x-3}\]

OpenStudy (anonymous):

\[\sqrt{x+3}\], sory its plus

OpenStudy (ranga):

sqrt(x+3) is only defined for x >= -3. It is a function that ALWAYS increases as x increases. Therefore, can you guess where it will have a minimum?

OpenStudy (anonymous):

i dont get the second line

OpenStudy (ranga):

Put x = -3, -2, -1, 0, 1, 2, 3, etc. and you can see the value of sqrt(x+3) will keep on increasing. That is what I meant in the second line.

OpenStudy (anonymous):

so would this be the case for every equation with a square root ?

OpenStudy (ranga):

It depends on the expression inside the square root. Here, inside the square root, you have x+3 which keeps on increasing as x increases and therefore sqrt(x+3) will also keep on increasing. But if you have some other expression, then it may increase and then decrease, etc. and so sqrt will also behave accordingly.

OpenStudy (anonymous):

like what if it was x-3 ?

OpenStudy (ranga):

x -3 is also a function that keeps on increasing as x increases. If you plot y = x - 3, you will see it is a straight line with a positive slope meaning it keeps on increasing as x increases. So sqrt(x-3) will also keep on increasing.

OpenStudy (ranga):

sqrt(1000 - x), for example, will keep on DECREASING as x increases.

OpenStudy (ranga):

The function, x^2 - 8x + 30 keeps decreasing for x < 4 and keeps increasing for x > 4. So sqrt(x^2 - 8x + 30) will also keep decreasing for x < 4 and keeps increasing for x > 4.

OpenStudy (anonymous):

so what would the solution be to the promblem ?

OpenStudy (anonymous):

a0 When x = -3 then (x - 3) = 0 b) when x = +infinity. If x = -infinity, the function is undefined.